# Add-One-In: Incremental Sample Selection for Large Language Models via a Choice-Based Greedy Paradigm

### v1

@@ Abstract
Selecting high-quality and diverse training samples from extensive datasets plays a crucial role in reducing training overhead and enhancing the performance of Large Language Models (LLMs). However, existing studies fall short in assessing the overall value of selected data, focusing primarily on individual quality, and struggle to strike an effective balance between ensuring diversity and minimizing data point traversals. Therefore, this paper introduces a novel choice-based sample selection framework that shifts the focus from evaluating individual sample quality to comparing the contribution value of different samples when incorporated into the subset. Thanks to the advanced language understanding capabilities of LLMs, we utilize LLMs to evaluate the value of each option during the selection process. Furthermore, we design a greedy sampling process where samples are incrementally added to the subset, thereby improving efficiency by eliminating the need for exhaustive traversal of the entire dataset with the limited budget. Extensive experiments demonstrate that selected data from our method not only surpass the performance of the full dataset but also achieves competitive results with state-of-the-art (SOTA) studies, while requiring fewer selections. Moreover, we validate our approach on a larger medical dataset, highlighting its practical applicability in real-world applications. [^1] https://github.com/BIRlz/comperative_sample_selection.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional success in Natural Language Processing (NLP) Brown et al. (2020); Touvron et al. (2023); OpenAI (2024); Du et al. (2025). Following the knowledge-based pre-training stage, human-oriented supervised instruction fine-tuning (SFT) Wei et al. (2022) significantly improves LLMs with the most increased performance. However, huge number of parameters and high complexity of these models also lead to substantial computational and financial demands during SFT, especially when faced with extensive training data.
Recently, some studies have indicated that not all instruction data equally contribute to fine-tuning, with a most representative subset often sufficient to match or surpass the performance of LLMs tuned on full datasets Zhou et al. (2023). Moreover, many work have shown that the diversity and quality of SFT data are crucial for fully unleashing the potential of LLM Ouyang et al. (2022); Xu et al. (2023); Du et al. (2024); Li et al. (2024d). Therefore, there is a growing interest in developing sample selectors to identify optimal subsets for more efficient SFT. The construction of data selectors relies on the design of selection criteria, considering both the sources of quality labels and approaches to obtain them, which form the fundamental difference for judging data quality. As highlighted by  Liu et al. (2024b), current approaches broadly adopt two strategies. The first one leverages sample internal information and solely relies on its intrinsic characteristics (Li et al., 2024c). The second strategy incorporates quality labels derived from external sources such as LLM preference judgments (Chen et al., 2024; Liu et al., 2024a) or continuous influence scores quantifying a sample’s impact on model behavior (Xia et al., 2024; Cao et al., 2024).
@@ 2 Related Work
The field of instruction sample selection for LLMs has seen significant advancements in recent years, driven by the need to improve model performance while reducing training costs Liu et al. (2024b). Early efforts focus on general data selection methods, often relying on human-designed features or simple heuristics to identify high-quality and more-diverse samples. For instance, Instruction-Mining Cao et al. (2024) and InstructionGPT-4 Wei et al. (2023b) utilize linguistic indicators and GPT-4 scores to guide the selection process, aiming to capture the quality of data through surface-level features. However, these methods often fell short in capturing the nuanced interactions between data and model performance. A more targeted approach emerged with the introduction of model- and data-centric selection criteria, where methods like IFD Li et al. (2024c) and SuperFiltering Li et al. (2024b) leverage internal information from the candidate dataset and the target model itself to select data, such as instruction following difficulty and perplexity of each sample. However, these methods are highly model-dependent and rely on the pre-experience training of the target LLM.
@@
In addition, there are also various selection methods that target on facilitating LLMs’ pre-training, like D4 Tirumala et al. (2023), DSIR Xie et al. (2023), and QuRating Wettig et al. (2024). Besides, Agrawal et al. (2022) and Nguyen and Wong (2023) expand sample selection into in-context learning and machine translation. However, we focus on empowering LLMs with stronger instruction-following capability in SFT.
@@ 3 Method
[…]
@@ 3.1 Warming-Up
3.2 A LLM-Driven Greedy Method For Contribution-Oriented Subset Selection
To efficiently find a desirable subset $\mathcal{A}$ without incurring the computational overhead of evaluating each candidate individually, we propose a LLM-driven greedy method that leverages the expressive capabilities of LLMs in understanding of the input and capable of handling the long-context to select the most valuable sample in a single computation at each iteration. Specifically, in each selection iteration, we firstly fill a selection prompt with $\mathcal{A}$ and $\mathcal{B}$ simultaneously, then query a LLM to make a choice about which element in $\mathcal{B}$ exhibits higher quality and can contribute the most diversity to $\mathcal{A}$.
As shown in the Fig. 2, we reformulate the selection obj. 1 by serving a LLM as the implementation of $F$ by such a carefully designed selection prompt. In order to prevent the LLM from failing to understand this complex selection prompt well, resulting in mis-following and ultimately being unable to make effective decisions, we adopt an ICL strategy Dong et al. (2022) to assist the LLM in better accomplishing our selection task and finally obtaining an effective optimal element. Finally, we can obtain an optimal element $x^{*}$ from the candidate set $\mathcal{B}$ by prompting $F_{\text{LLM}}$.
@@ 3.2 A LLM-Driven Greedy Method For Contribution-Oriented Subset Selection
where sizes of $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ are denoted as $L_{A}$ and $L_{B}$, respectively. Our LLM-driven greedy selection process are summarized in App. A, where our method begins with a random initialization of $\mathcal{A}$ and without specific statement, we set $L_{A}=L_{B}=20$.
@@ 3.3 Discussion
Moreover, our method parallels the classical greedy algorithm used in submodular optimization. when a LLM can correctly follow the selection prompt, our greedy algorithm could achieve at least a $(1-{e}^{-\gamma})$ approximation of the optimal solution (Chen et al., 2018), where the parameter $\gamma\in[0,1)$ is a submodularity ratio related to the LLM. This means that our method not only efficiently builds a diverse and high-quality subset but also provides theoretical assurance on the solution’s near-optimality. In terms of computational cost, each iteration involves querying the LLM to select the next sample, resulting in a total time complexity of $\mathcal{O}(K\cdot T_{\text{LLM}})$, where $K$ is the desired subset size and $T_{\text{LLM}}$ is the inference time of the LLM. This linear complexity with respect to $K$ makes our method scalable to large datasets and requires fewer accesses to the entire dataset.
@@ 4.1.1 Settings
In order to comprehensively evaluate the flexibility of our method, we employ three types of LLM as the selector: 1) Small-size: Llama3.1-8B-Instruct Touvron et al. (2023); 2) Medium-size: Qwen2.5-72B-Instruct Yang et al. (2024); 3) Closed-source: ChatGPT / GPT4o OpenAI (2024).
@@ 4.1.3 Results
By following AlpaGasus, which selects the 9K highest-quality samples scored by ChatGPT, we first demonstrate the effectiveness of our method using the same scale of selected samples. As shown in Fig. 3, our model trained with 9K samples significantly outperforms the model trained on the full dataset under various settings. We employ a diverse range of LLMs as selectors and as models to be trained (Llama2-7B-hf and Llama2-13B-hf). These results illustrate that our method effectively identifies valuable and diverse samples for instruction fine-tuning, leveraging the powerful language understanding capability of LLMs, regardless of the specific size or family of models used for selection or training.
Furthermore, we conduct experiments by selecting subsets corresponding to different percentages of the training dataset, ranging from 5% to 20%, and compare the average win score changes relative to the full dataset as the amount of data increases. WS can be calculated as #$\frac{\#\text{Win}-\#\text{Lose}}{\#\text{All}}+1$, where a score greater than 1.0 indicates that the model outperforms the one that is compared against. As shown in Fig. 4, with just 5% of the data it selects, our model can outperform those trained on the full dataset on three of the four test datasets. As the amount of data increases, our method consistently exceeds the performance of the Full across all four datasets. In particular, with 15% data usage, our model achieves the best performance, with an average win rate of 1.257. These results demonstrate that appropriately selecting a subset of the full data is sufficient to enhance the LLM’s instruction-following ability, and our incremental method effectively identifies a high-quality and more diverse subset for improved LLM tuning.
@@
In our comparison with IFD, our method demonstrates competitive performance, as illustrated by the red dotted line in Fig.4. We consistently outperform IFD across various percentages ranging from 5% to 20%, achieving an optimal win score of 1.17 at 15%, which clearly highlights the strength and practical utility of our selection approach. Using the powerful understanding capabilities of LLMs, we can incrementally identify a diverse subset without having to examine the entire dataset at once, thereby greatly improving efficiency by reducing the number of selection times. Additionally, we provide comparisons with the Random baseline in App.B.2, where our method also achieves a significant performance improvement.
@@
Following practice of evaluating LLMs on benchmarks, we also compare our method with Full, Random, AlpaGasus, and IFD, where we set sample size to 9K for a fair comparison, except for Full. For MMLU, BBH, and Hellaswag, we adopt 0/5-Shot, 3-Shot, and 0-Shot settings, respectively. Tab. 1 suggests that our method provides a promising approach to diverse data selection by achieving competitive performance to baselines with fewer selection times, optimizing both performance and selection cost.
@@ 4.2 Analysis
In this section, we conduct an in-depth analysis of our method from the perspective of hyperparameter sensitive, efficiency analysis, characteristics of the selected samples, and case studies.
@@
As summarized in Fig. 6, we compare different configurations in terms of their ability to select better samples and report pairwise comparison results, indicating superiority in diversity and quality of the selected subsets. Our key findings are as follows: 1) As shown in Fig. 6(1), (2), and (3), increased $L_{A}$ or $L_{B}$ leads to better performance consistently across different datasets and model selectors. A larger $L$ provides the LLMs with more context and a broader range of candidates to choose from, which in turn enhances the diversity and quality of the selected samples. 2) Fig. 6(4) illustrates that our method effectively integrates with a multi-round selection strategy. By adopting a hierarchical, coarse-to-fine approach, we iteratively refine the subset through multiple selection rounds, gradually narrowing the focus to increasingly promising candidates. The data selected in each round serves as the foundation for constructing new options in the subsequent round, enabling a step-by-step process to identify higher-quality data over iterations. Additional results based on ChatGPT-as-the-selector are in App. B.3.
@@
We investigate whether our method achieves an acceptable trade-off between improved performance and the computational costs associated with the selection, from the perspective of: 1) Time Complexity; 2) Number of Iterations; 3) Practical Time Consumption. Using the Alpaca Dataset (52K), we present an efficiency comparison when selecting the top 10% of the data. The results in Tab. 2 demonstrate that our method significantly reduces selection costs compared to IFD and AlpaGasus. Specifically, operating with a time complexity proportional to $K$ rather than $N$, our method requires significantly fewer iterations and less practical time, due to the greedy selection process that incorporates the powerful capability of LLMs. This not only accelerates the selection procedure, but also maintains high performance, achieving better efficiency without compromising the quality of the selected data. In summary, our method is only proportional to the desired subset size $K$, embodying the principle of “selecting only what is necessary”. This characteristic will ensure that our method can guarantee higher efficiency and lower selection costs when faced with extremely large-scale datasets. We provide implementation details in App. B.5.
@@
We evaluate whether our method identifies a representative subset with higher quality and greater diversity. Tab. 3 presents a comparison using various diversity metrics, such as Type-Token Ratio (TTR) Richards (1987), Measure of Textual Lexical Diversity (MTLD) McCarthy and Jarvis (2010), and Simpson Diversity Index (SDI) Simpson (1997), along with quality metrics including DEITA quality score Liu et al. (2024a) and helpful reward value Dong et al. (2023). Higher values generally indicate better diversity or quality, except for the SDI. We also report instruction token counts to assess structural similarity to the full dataset. Baseline methods are evaluated on 9K samples versus 52K with average scores reported. We provide implementation details and explanations to these metrics in App. B.4.
@@
Please refer to App. C.

### v2

@@ Abstract
Selecting high-quality and diverse training samples from extensive datasets plays a crucial role in reducing training overhead and enhancing the performance of Large Language Models (LLMs). However, existing studies fall short in assessing the overall value of selected data, focusing primarily on individual quality, and struggle to strike an effective balance between ensuring diversity and minimizing data point traversals. Therefore, this paper introduces a novel choice-based sample selection framework that shifts the focus from evaluating individual sample quality to comparing the contribution value of different samples when incorporated into the subset. Thanks to the advanced language understanding capabilities of LLMs, we utilize LLMs to evaluate the value of each option during the selection process. Furthermore, we design a greedy sampling process where samples are incrementally added to the subset, thereby improving efficiency by eliminating the need for exhaustive traversal of the entire dataset with the limited budget. Extensive experiments demonstrate that selected data from our method not only surpasses the performance of the full dataset but also achieves competitive results with recent powerful studies, while requiring fewer selections. Moreover, we validate our approach on a larger medical dataset, highlighting its practical applicability in real-world applications. Our code and data are available at https://github.com/BIRlz/comperative_sample_selection.
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); OpenAI (2024) have demonstrated exceptional success in AI Chiang et al. (2023); Hu et al. (2025b); Li et al. (2025); Dai et al. (2025); Hu et al. (2023, 2022, 2025a, 2024). Following the knowledge-based pre-training stage, human-oriented supervised fine-tuning (SFT) Wei et al. (2022); Du et al. (2025) significantly improves LLMs with the most increased performance. However, a huge number of parameters and high complexity of these models also lead to substantial computational and financial demands during SFT, especially when faced with extensive training data.
Recently, some studies have indicated that not all instruction data equally contribute to fine-tuning, with a most representative subset often sufficient to match or surpass the performance of LLMs tuned on full datasets Zhou et al. (2023). Moreover, many work have shown that the diversity and quality of SFT data are crucial for fully unleashing the potential of LLMs Ouyang et al. (2022); Xu et al. (2023). Therefore, there is a growing interest in developing sample selectors to identify optimal subsets for more efficient SFT. The construction of data selectors relies on the design of selection criteria, considering both the sources of quality labels and approaches to obtain them, which form the fundamental difference for judging data quality. As highlighted by  Liu et al. (2024b), current approaches broadly adopt two strategies. The first one leverages sample internal information and solely relies on its intrinsic characteristics (Li et al., 2024c). The second strategy incorporates quality labels derived from external sources such as LLM preference judgments (Chen et al., 2024; Liu et al., 2024a) or continuous influence scores quantifying a sample’s impact on model behavior (Xia et al., 2024; Cao et al., 2024).
@@ 2 Related Work
The field of instruction sample selection for LLMs has seen significant advancements in recent years, driven by the need to improve model performance (i.e., safety Du et al. (2024)) while reducing training costs Liu et al. (2024b). Early efforts focus on general data selection methods, often relying on human-designed features or simple heuristics to identify high-quality and more diverse samples. For instance, Instruction-Mining Cao et al. (2024) and InstructionGPT-4 Wei et al. (2023b) utilize linguistic indicators and GPT-4 scores to guide the selection process, aiming to capture the quality of data through surface-level features. However, these methods often fell short in capturing the nuanced interactions between data and model performance. A more targeted approach emerged with the introduction of model- and data-centric selection criteria, where methods like IFD Li et al. (2024c) and SuperFiltering Li et al. (2024b) leverage internal information from the candidate dataset and the target model itself to select data, such as instruction following difficulty and perplexity of each sample. However, these methods are highly model-dependent and rely on the pre-experience training of the target LLM.
@@
In addition, there are also various selection methods that target on facilitating LLMs’ pre-training, like D4 Tirumala et al. (2023), DSIR Xie et al. (2023), and QuRating Wettig et al. (2024). Besides, Agrawal et al. (2022) and Nguyen and Wong (2023) expand sample selection into in-context learning and machine translation, while we focus on SFT.
@@ 3 Method
Let $\mathcal{D}=\{x_{1},x_{2},\ldots,x_{N}\}$ denote an instruction tuning dataset with size $N$. Usually each sample $x_{i}$ is a triplet $\{\text{Instruction},[\text{Input}],\text{Answer}\}$, where $\text{Input}]$ is an optional part associated with the instruction. Given a trainable LLM parameterized by $\theta\in\mathbb{R}^{d}$, we denote $\theta_{\mathcal{D}}$ as the instruction fine-tuned LLM $\theta$ on dataset $\mathcal{D}$. Our objective is to effectively find a subset $\mathcal{A}\subseteq\mathcal{D}$ ($|\mathcal{A}|=K\ll|\mathcal{D}|$) with $K$ data budget, such that each selected sample in $\mathcal{A}$ satisfies specific criteria defined by a selection function $F(\cdot)$. Moreover, $\theta_{\mathcal{A}}$ can achieve comparable performance than $\theta_{\mathcal{D}}$ on various downstream tasks.
@@ 3.1 Warming-Up
3.2 An LLM-Driven Greedy Method For Contribution-Oriented Subset Selection
To efficiently find a desirable subset $\mathcal{A}$ without incurring the computational overhead of evaluating each candidate individually, we propose an LLM-driven greedy method that leverages the expressive capabilities of LLMs in understanding of the input and capable of handling the long-context to select the most valuable sample in a single computation at each iteration. Specifically, in each selection iteration, we firstly fill a selection prompt with $\mathcal{A}$ and $\mathcal{B}$ simultaneously, then query an LLM to make a choice about which element in $\mathcal{B}$ exhibits higher quality and can contribute the most diversity to $\mathcal{A}$.
As shown in the Fig. 2, we reformulate the selection obj. 1 by serving an LLM as the implementation of $F$ by such a carefully designed selection prompt. In order to prevent the LLM from failing to understand this complex selection prompt well, resulting in mis-following and ultimately being unable to make effective decisions, we adopt an ICL strategy Dong et al. (2022) to assist the LLM in better accomplishing our selection task and finally obtaining an effective optimal element. Finally, we can obtain an optimal element $x^{*}$ from the candidate set $\mathcal{B}$ by prompting $F_{\text{LLM}}$.
@@ 3.2 An LLM-Driven Greedy Method For Contribution-Oriented Subset Selection
where sizes of $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ are denoted as $L_{A}$ and $L_{B}$, respectively. Our LLM-driven greedy selection process are summarized in App. 1, where our method begins with a random initialization of $\mathcal{A}$ and without specific statement, we set $L_{A}=L_{B}=20$.
Input: $\mathcal{D}$: Full dataset, an LLM and hyper-parameters: {$K$: Desired subset size, $L_{A}$ and $L_{B}$: window size.}
Output: $\mathcal{A}$: Selected subset
3
4while $|\mathcal{A}|<K$ do
6
8 update $\mathcal{A}\leftarrow\mathcal{A}\cup\{x^{*}\}$;
9 update $\mathcal{B}\leftarrow\mathcal{B}\setminus\{x^{*}\}$;
10
return $\mathcal{A}$
Algorithm 1 Algorithm Process of Our Method.
@@ 3.3 Discussion
Moreover, our method parallels the classical greedy algorithm used in submodular optimization. when an LLM can correctly follow the selection prompt, our greedy algorithm could achieve at least a $(1-{e}^{-\gamma})$ approximation of the optimal solution (Chen et al., 2018), where the parameter $\gamma\in[0,1)$ is a submodularity ratio related to the LLM. This means that our method not only efficiently builds a diverse and high-quality subset but also provides theoretical assurance on the solution’s near-optimality. In terms of computational cost, each iteration involves querying the LLM to select the next sample, resulting in a total time complexity of $\mathcal{O}(K\cdot T_{\text{LLM}})$, where $K$ is the desired subset size and $T_{\text{LLM}}$ is the inference time of the LLM. This linear complexity with respect to $K$ makes our method scalable to large datasets and requires fewer accesses to the entire dataset.
@@ 4.1.1 Settings
In order to comprehensively evaluate the flexibility of our method, we employ three types of LLM as the selector: 1) Small-size: Llama3.1-8B-Instruct Touvron et al. (2023); 2) Medium-size: Qwen2.5-72B-Instruct Yang et al. (2024); 3) Closed-source: GPT3.5 / GPT4o OpenAI (2024).
@@ 4.1.3 Results
By following AlpaGasus, which selects the 9K highest-quality samples scored by GPT3.5, we first demonstrate the effectiveness of our method using the same scale of selected samples. As shown in Fig. 3, our model trained with 9K samples significantly outperforms the model trained on the full dataset under various settings. We employ a diverse range of LLMs as selectors and as models to be trained (Llama2-7B-hf and Llama2-13B-hf). These results illustrate that our method effectively identifies valuable and diverse samples for instruction fine-tuning, leveraging the powerful language understanding capability of LLMs, regardless of the specific size or family of models used for selection or training.
Furthermore, we conduct experiments by selecting subsets corresponding to different percentages of the training dataset, ranging from 5% to 20%, and compare the average win score changes relative to the full dataset as the amount of data increases. WS can be calculated as $\frac{\#\text{Win}-\#\text{Lose}}{\#\text{All}}+1$, where a score greater than 1.0 indicates that the model outperforms the one that is compared against. As shown in Fig. 4, with just 5% of the data it selects, our model can outperform those trained on the full dataset on three of the four test datasets. As the amount of data increases, our method consistently exceeds the performance of the Full across all four datasets. In particular, with 15% data usage, our model achieves the best performance, with an average win rate of 1.257. These results demonstrate that appropriately selecting a subset of the full data is sufficient to enhance the LLM’s instruction-following ability, and our incremental method effectively identifies a high-quality and more diverse subset for improved LLM tuning.
@@
In our comparison with IFD, our method demonstrates competitive performance, as illustrated by the red dotted line in Fig.4. We consistently outperform IFD across various percentages ranging from 5% to 20%, achieving an optimal win score of 1.17 at 15%, which clearly highlights the strength and practical utility of our selection approach. Using the powerful understanding capabilities of LLMs, we can incrementally identify a diverse subset without having to examine the entire dataset at once, thereby greatly improving efficiency by reducing the number of selection times. Additionally, we provide comparisons with the Random baseline in Fig.6, where our method consistently outperforms the Random baseline across various settings by significant margins, demonstrating the effectiveness of the proposed method.
@@
Following the practice of evaluating LLMs on benchmarks, we also compare our method with Full, Random, AlpaGasus, and IFD, where we set the sample size to 9K for a fair comparison, except for Full. For MMLU, BBH, and Hellaswag, we adopt 0/5-Shot, 3-Shot, and 0-Shot settings, respectively. Tab. 1 suggests that our method provides a promising approach to diverse data selection by achieving competitive performance to baselines with fewer selection times, optimizing both performance and selection cost. More detailed performance with varying scales of training data can be found in App. A.2.
@@
Table 1: The benchmark results of models fine-tuned on different subsets selected by corresponding methods. More performance and analysis with varying training samples can be found in Tab. 7.
@@ 4.2 Analysis
In this section, we conduct an in-depth analysis of our method from the perspective of hyperparameter sensitivity, efficiency analysis, characteristics of the selected samples, and case studies.
@@
Fig. 7 compares different configurations in terms of their ability to select better samples and report pairwise comparison results, indicating superiority in diversity and quality of the selected subsets. Our key findings are as follows: 1) As shown in Fig. 7(1), (2), and (3), increased $L_{A}$ or $L_{B}$ leads to better performance consistently across different datasets and model selectors. A larger $L$ provides the LLMs with more context and a broader range of candidates to choose from, which in turn enhances the diversity and quality of the selected samples. 2) Fig. 7(4) illustrates that our method effectively integrates with a multi-round selection strategy. By adopting a hierarchical, coarse-to-fine approach, we iteratively refine the subset through multiple selection rounds, gradually narrowing the focus to increasingly promising candidates. The data selected in each round serves as the foundation for constructing new options in the subsequent round, enabling a step-by-step process to identify higher-quality data over iterations. Additional results based on GPT3.5-as-the-selector are in App. A.3.
@@
We investigate whether our method achieves an acceptable trade-off between improved performance and the computational costs associated with the selection, from the perspective of: 1) Time Complexity; 2) Number of Iterations; 3) Practical Time Consumption. Using the Alpaca Dataset (52K), we present an efficiency comparison when selecting the top 10% of the data. The results in Tab. 2 demonstrate that our method significantly reduces selection costs compared to IFD and AlpaGasus. Specifically, operating with a time complexity proportional to $K$ rather than $N$, our method requires significantly fewer iterations and less practical time, due to the greedy selection process that incorporates the powerful capability of LLMs. This not only accelerates the selection procedure, but also maintains high performance, achieving better efficiency without compromising the quality of the selected data. In summary, our method is only proportional to the desired subset size $K$, embodying the principle of “selecting only what is necessary”. This characteristic will ensure that our method can guarantee higher efficiency and lower selection costs when faced with extremely large-scale datasets. We provide implementation details in App. A.5.
@@
We evaluate whether our method identifies a representative subset with higher quality and greater diversity. Tab. 3 presents a comparison using various diversity metrics, such as Type-Token Ratio (TTR) Richards (1987), Measure of Textual Lexical Diversity (MTLD) McCarthy and Jarvis (2010), and Simpson Diversity Index (SDI) Simpson (1997), along with quality metrics including DEITA quality score Liu et al. (2024a) and helpful reward value Dong et al. (2023). Higher values generally indicate better diversity or quality, except for the SDI. We also report instruction token counts to assess structural similarity to the full dataset. Baseline methods are evaluated on 9K samples versus 52K with average scores reported. We provide implementation details and explanations to these metrics in App. A.4.
@@
We highlight three selection cases in App. B.